# Animate Your Thoughts: Decoupled Reconstruction of Dynamic Natural Vision from Slow Brain Activity

### v1

@@ 6 Conclusion
In this paper, we introduce a video reconstruction model (Mind-Animator) that decouples semantic, structural, and motion information from fMRI, achieving state-of-the-art performance across 3 public datasets. We mitigate the interference of external video data on motion information decoding through a rational experimental design. The results of the permutation test demonstrate that the motion information we decoded indeed originates from fMRI, rather than being a "hallucination" from generative model. Additionally, the visualization of voxel-wise and ROI-wise importance maps substantiate the neurobiological interpretability of our model.
Acknowledgments and Disclosure of Funding
We would like to express our gratitude to Prof.Jack L. Gallant and Prof.Shinji Nishimoto for their pioneering exploration in the field of video reconstruction and for their high-quality code. We are grateful to Prof.Juan Helen Zhou and Dr.Zijiao Chen for their patient answers to our questions and for making all the results of the Mind-video test set public. We also extend our thanks to Prof.Michal Irani, Dr.Ganit Kupershmidt, and Dr.Roman Beliy for providing us with all the reconstruction results of their models on the test set.
We would like to express our appreciation to Prof.Zhongming Liu and Dr.Haiguang Wen for their open-sourced high-quality video-fMRI dataset and the preprocessing procedures. Our gratitude also goes to the Human Connectome Project (HCP) for providing a large-scale fMRI dataset and cortical visualization tools. We are thankful to the Algonauts2021 competition for providing a set of pre-processed video-fMRI data from multiple subjects.
We are thankful to the Stable Diffusion team for their high-performance text-to-image model, and we also appreciate the Tune-a-video team for their open-source video editing framework, which allows us to reconstruct videos without introducing additional motion information.
@@ A.1 Reconstructing Human Vision from Brain Activities
Building on Haxby’s[60] seminal work, the field of neural decoding has seen a proliferation of tasks with significant implications for guiding research. These tasks can be broadly classified into three categories: stimulus classification, identification, and reconstruction, with the latter being the most challenging and the focus of our study.
Traditional image reconstruction techniques rely on linear regression models to correlate fMRI with manually defined image features[61, 12, 8], yielding blurry results and a heavy reliance on manual feature selection. However, the advent of deep learning has revolutionized this domain. Deep neural networks (DNNs) have become increasingly prevalent for their ability to address the scarcity of stimulus-fMRI pairs through semi-supervised learning[62], as demonstrated by Beliy et al.[22] and Gaziv et al.[63]. Yet, these models often fail to capture discernible semantic information. Chen et al.[19] employed a pre-training and fine-tuning approach on fMRI data, leveraging methods akin to Masked Autoencoder(MAE)[64] and Latent Diffusion Models(LDM)[38] to improve reconstruction quality. Ozcelik et al.[21] and Gu et al.[65] utilized self-supervised models for feature extraction, followed by iterative optimization to refine the reconstruction process. The integration of semantic information from text, facilitated by Contrastive Language-Image Pre-Training (CLIP)[37] , has been instrumental in reconstructing complex natural images. Lin et al.[16] and Takagi et al.[20] demonstrated the potential of aligning fMRI with CLIP representations and mapping fMRI to text and image features for high-fidelity reconstruction.
@@
Compared to image reconstruction, the challenge in video reconstruction lies in the significant discrepancy between the temporal resolution of fMRI (0.5Hz) and the frame rate of the stimulus video (30Hz), which presents a substantial challenge in modeling the mapping between fMRI signals and video content. To overcome the challenge, Nishimoto [28] transformed the video reconstruction task into a identification problem, employing the Motion-Energy model [29] and Bayesian inference to reconstruct videos from a predefined video library. Subsequently, Han [30] and Wen et al.[31] mapped brain responses to the feature spaces of DNN to reconstruct down-sampled (with the frame rate reduced to 1Hz) video stimuli. The aforementioned studies have preliminarily validated the feasibility of reconstructing video frames from fMRI. Wang et al.[32] developed an f-CVGAN that learns temporal and spatial information in fMRI through separate discriminators [33]. To mitigate the scarcity of fMRI-video data, Kupershmidt et al.[34] utilized self-supervised learning [35] to incorporate a large amount of unpaired video data. These efforts have validated the feasibility of video reconstruction from fMRI, albeit with a lack of explicit semantic information in the results. Chen et al.[36] utilized contrastive learning to map fMRI to the CLIP representation space and fine-tuned inflated Stable Diffusion [38, 39] on a video-text dataset as a video generation model, successfully reconstructing coherent videos with clear semantic information for the first time. However, Chen did not consider structure information such as color and position, and it was uncertain whether the motion information in the reconstructed videos originated from the fMRI or the video generation model.
@@ A.2 Diffusion Models
Diffusion models[66, 67] , a class of probabilistic generative models, have increasingly rivaled or surpassed the performance of Generative Adversarial Networks (GAN)[33] in specific tasks within the field of computer vision. Diffusion models encompass a forward diffusion process and a reverse denoising process, each exhibiting Markovian behavior. The forward process incrementally introduces Gaussian noise into the original image, culminating in a transition to standard Gaussian noise. The forward diffusion process can be represented as $q(x_{t}|x_{t-1})=N(x_{t};\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I)$, where $t$ denotes the time step of each noise addition. The reverse denoising process employs the U-Net[68] architecture to accurately model the noise distribution at each timestep $t$. The image synthesis is achieved through a sequential denoising and sampling procedure, initiated from standard Gaussian noise.
In the context of image generation tasks, the conventional diffusion model executes two Markov processes in a large pixel space, resulting in substantial computational resource utilization. To address this issue, Latent Diffusion Models (LDM)[38] employs a VQ-VAE[40] encoder to transform the pixel space into a low-dimensional latent space. Subsequently, the diffusion model’s training and generation are performed in the latent space, with the final generated image obtained by utilizing the VQ-VAE decoder. This approach significantly reduces computational resource requirements and inference time while preserving the quality of generated images.
@@ A.3 Diffusion Models for Video Generation
After achieving significant progress in text-to-image (T2I) generation tasks, diffusion models have piqued the interest of researchers in exploring their potential for text-to-video (T2V) generation. The pioneering work by Ho et al.[69] , introducing the 3D diffusion U-Net, marked significant progress in applying Diffusion Models to video generation. This was followed by further advancements by Ho et al.[70] , who utilized a cascaded sampling framework and super-resolution method to generate high-resolution videos. Subsequent contributions have expanded upon this work, notably with the incorporation of a temporal attention mechanism over frames by Singer et al.[71] in Make-A-Video. Zhou et al.[72] with MagicVideo, and He et al.[73] with LVDM, have integrated this mechanism into latent Diffusion Models, significantly enhancing video generation capabilities. However, due to the scarcity of paired text-video datasets and the high memory requirements for training 3D U-Nets, alternative approaches are being explored. These involve refining pre-trained T2I models to directly undertake T2V tasks. Khachatryan et al.[47] introduced two enhancements to enable zero-shot adaptation of T2I models to T2V tasks: (1) the implementation of cross-frame attention, ensuring that the generation of each current frame in a video considers information from preceding frames; and (2) the consideration of inter-frame correlations during noise sampling, rather than random sampling for each frame independently. Wu et al.[39] also employed cross-frame attention and achieved one-shot video editing by fine-tuning partial model parameters on individual videos.
@@
Appendix B Model Architecture
B.1 Consistency Motion Generator
The Temporal Module is tasked with learning the temporal dynamics from the visible frames. Given the severe information redundancy between video frame tokens, we specifically design a Sparse Causal mask. As shown in Figure 3 on the right, during training, the mask is divided into fixed and random components. The fixed mask ensures that each frame cannot access information from subsequent frames, while the random mask maintains sparsity among visible frames, preventing the model from taking shortcuts[52] and accelerating training. During inference, we eliminate the random mask to allow full utilization of information from all preceding frames for predicting future frames.
B.2 Text-to-Image Network Inflation
To leverage pre-trained weights from large-scale image datasets, such as ImageNet, for the pre-training of massive video understanding models, Carreira et al.[46] pioneered the expansion of filters and pooling kernels in 2D ConvNets into the third dimension to create 3D filters and pooling kernels. This process transforms N×N filters used for images into N×N×N 3D filters, providing a beneficial starting point for 3D video understanding models by utilizing spatial features learned from large-scale image datasets.
Appendix C Data Preprocessing
For the stimulus videos of the three datasets described below, we segmented them into 2-second clips, down-sampled the frame rate to 4Hz (i.e., evenly extracting 8 frames), then centrally cropped each frame, and resized each to a shape of 512x512. Following the approach of Chen et al.[36], we employed BLIP2[74] to obtain textual descriptions for each video clip, with lengths not exceeding 20 words.
C.1 Video Captioning with BLIP2
PyTorch code for the video captioning process is depicted in Algorithm 1.
@@ C.1 Video Captioning with BLIP2
import torch
import clip
import random
import numpy as np
from PIL import Image
from lavis.models import load_model_and_preprocess
device = torch.device(’cuda:5’)
clip_model, clip_preprocess = clip.load("ViT-B/32", device=device)
model,vis_processors,_=load_model_and_preprocess(name="blip2_t5",
model_type="pretrain_flant5xxl", is_eval=True, device=device)
prompt = "Question:Whatdoesthisimagedescribe?Answer:"
def Video_Captioning(Train_video_path_root):
Train_Captions = []
for i in tqdm(range(18)):
for j in tqdm(range(240)):
@@
frame1 = Image.open(frames_root + ’0000000.jpg’).convert("RGB")
frame1 = vis_processors["eval"](frame1).unsqueeze(0).to(device)
frame2 = Image.open(frames_root + ’0000056.jpg’).convert("RGB")
frame2 = vis_processors["eval"](frame2).unsqueeze(0).to(device)
frame_mid = Image.open(frames_root + ’0000024.jpg’).convert("RGB")
clip_image = clip_preprocess(frame_mid).unsqueeze(0).to(device)
caption1 = model.generate({"image": frame1, "prompt": prompt})
caption2 = model.generate({"image": frame2, "prompt": prompt})
text1 = clip.tokenize(caption1).to(device)
text2 = clip.tokenize(caption2).to(device)
with torch.no_grad():
image_features3 = clip_model.encode_image(clip_image)
text_features1 = clip_model.encode_text(text1)
text_features2 = clip_model.encode_text(text2)
cos_sim1 = torch.cosine_similarity(image_features3, text_features1)
cos_sim2 = torch.cosine_similarity(image_features3, text_features2)
if abs(cos_sim1 - cos_sim2) <= 0.05:
number = random.random()
if number >= 0.5:
caption = caption1[0]
else:
caption = caption2[0]
else:
caption = caption1[0] + ’,andthen + caption2[0]
Train_Captions.append(caption)
Train_Captions = np.array(Train_Captions)
return Train_Captions
C.2 CC2017
@@
CC2017[31] dataset was first used in the work of Wen et al.[31] This dataset include fMRI data from 3 subjects who view a variety of movie clips (23° × 23°) with a central fixation cross (0.8° × 0.8°). Clips are divided into 18 training movies and 5 testing movies, each eight minutes long, presented 2 and 10 times to each subject, respectively. MRI (T1 and T2-weighted) and fMRI data (2-second temporal resolution) are acquired using a 3-T system. The fMRI volumes are processed for artifact removal, motion correction (6 DOF), registered to MNI space, and projected onto cortical surfaces coregistered to a template.
@@
To extract voxels in the activated visual areas, we calculate the correlation of the time series for each voxel’s activation across 2 trials within the training set. Subsequently, we apply Fisher’s z-transform to the computed correlations, average the results across 18 sessions, and identify the most significant 4500 voxels using a paired t-test to form a mask. This mask is computed separately for each of the 3 subjects on their respective training sets and then applied to both training and test set data, with the selected voxels averaged across trials. Following the work of Nishimoto et al.[28] and Han et al. [30] , we utilize BOLD signals with a 4-second lag to represent the movie stimulus responses, thereby accounting for the hemodynamic response delay.
C.3 HCP
@@
This dataset is part of the Human Connectome Project (HCP)[48], encompassing BOLD (blood-oxygen-level dependent) responses from 158 subjects. For the subsequent experiments, three subjects (100610, 102816, and 104416) are randomly selected from this dataset. Data acquisition is performed using a 7T MRI scanner with a spatial resolution of 1.6 millimeters and a repetition time (TR) of 1 second. The utilized BOLD signals undergo standard HCP preprocessing procedures, which include correction for head motion and distortion, high-pass filtering, and removal of temporal artifacts via independent component analysis (ICA). The preprocessed BOLD responses are then registered to the MNI standard space.
@@
Due to the difficulty in directly acquiring fMRI data from multiple trials, we directly utilize the parcellation of the human cerebral cortex proposed by Glasser et al.[75] to extract voxels within the activated visual cortex. The resulting ROIs we extract include: V1, V2, V3, hV4, PPA, FFA, LO, PHC, MT, MST, and TPOJ, totaling 5820 voxels. Following the work of Nishimoto et al.[28] and Han et al.[30] , we utilize BOLD signals with a 4-second lag to represent the movie stimulus responses, thereby accounting for the hemodynamic response delay. Given that no prior work has conducted video reconstruction experiments on these three subjects from the HCP dataset, except for Wang et al.[32], we randomly shuffle all video segments and allocate 90% for the training set, with the remaining 10% reserved for the test set.
C.4 Algonauts2021

### v2

@@ 7 Conclusion
We introduce a video reconstruction model (Mind-Animator) that decouples semantic, structural, and motion information from fMRI, achieving state-of-the-art performance across 3 public datasets. We mitigate the interference of external video data on motion information decoding through a rational experimental design. The results of the shuffle test demonstrate that the motion information we decoded indeed originates from fMRI, rather than being a spontaneity from generative model. Additionally, the visualization of voxel-wise and ROI-wise importance maps substantiate the neurobiological interpretability of our model.
8 Acknowledgement
This work was supported in part by the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0930000); in part by Beijing Natural Science Foundation under Grant L243016; and in part by the National Natural Science Foundation of China under Grant 62206284 and 62020106015.
@@ A.1 Reconstructing Human Vision from Brain Activities
Building on Haxby et al. (2001)’s seminal work, the field of neural decoding has seen a proliferation of tasks with significant implications for guiding research. These tasks can be broadly classified into three categories: stimulus classification, identification, and reconstruction, with the latter being the most challenging and the focus of our study.
Traditional image reconstruction techniques rely on linear regression models to correlate fMRI with manually defined image features (Kay (2008); Naselaris et al. (2009); Fujiwara et al. (2013)), yielding blurry results and a heavy reliance on manual feature selection. However, the advent of deep learning has revolutionized this domain. Deep neural networks (DNNs) have become increasingly prevalent for their ability to address the scarcity of stimulus-fMRI pairs through semi-supervised learning (Chapelle et al. (2009)), as demonstrated by Beliy et al. (2019) and Gaziv et al. (2022). Yet, these models often fail to capture discernible semantic information. Chen et al. (2023) employed a pre-training and fine-tuning approach on fMRI data, leveraging methods akin to Masked Autoencoder(MAE) (He et al. (2022a)) and Latent Diffusion Models(LDM) (Rombach et al. (2022)) to improve reconstruction quality. Ozcelik et al. (2022) and Gu et al. (2022) utilized self-supervised models for feature extraction, followed by iterative optimization to refine the reconstruction process. The integration of semantic information from text, facilitated by Contrastive Language-Image Pre-Training (CLIP) (Radford et al. (2021)), has been instrumental in reconstructing complex natural images. Lin et al. (2022) and Takagi & Nishimoto (2023) demonstrated the potential of aligning fMRI with CLIP representations and mapping fMRI to text and image features for high-fidelity reconstruction.
@@
Compared to image reconstruction, the challenge in video reconstruction lies in the significant discrepancy between the temporal resolution of fMRI (0.5Hz) and the frame rate of the stimulus video (30Hz), which presents a substantial challenge in modeling the mapping between fMRI signals and video content. To overcome the challenge, Nishimoto et al. (2011) transformed the video reconstruction task into a identification problem, employing the Motion-Energy model (Adelson & Bergen (1985)) and Bayesian inference to reconstruct videos from a predefined video library. Subsequently, Han et al. (2019) and Wen et al. (2018) mapped brain responses to the feature spaces of DNN to reconstruct down-sampled (with the frame rate reduced to 1Hz) video stimuli. Specifically, Han et al. mapped fMRI data to a VAEKingma & Welling (2014) pretrained on the ImageNet ILSVRC2012 Russakovsky et al. (2015) dataset to reconstruct a single frame, while Wen et al. mapped fMRI data to the feature space of AlexNetKrizhevsky et al. (2012) and used a deconvolutional neural networkZeiler et al. (2010) to reconstruct a single frame. The aforementioned studies have preliminarily validated the feasibility of reconstructing video frames from fMRI. Wang et al. (2022) developed an f-CVGAN that learns temporal and spatial information in fMRI through separate discriminators (Goodfellow et al. (2020)). To mitigate the scarcity of fMRI-video data, Kupershmidt et al. (2022) utilized self-supervised learning (Kingma & Welling (2014)) to incorporate a large amount of unpaired video data. These efforts have validated the feasibility of video reconstruction from fMRI, albeit with a lack of explicit semantic information in the results. Chen et al. (2024) utilized contrastive learning to map fMRI to the CLIP representation space and fine-tuned inflated Stable Diffusion (Rombach et al. (2022); Wu et al. (2023)) on a video-text dataset as a video generation model, successfully reconstructing coherent videos with clear semantic information for the first time. However, Chen did not consider structure information such as color and position, and it was uncertain whether the motion information in the reconstructed videos originated from the fMRI or the video generation model.
@@ A.2 Diffusion Models
Diffusion models (Wijmans & Baker (1995); Ho et al. (2020)), a class of probabilistic generative models, have increasingly rivaled or surpassed the performance of Generative Adversarial Networks (GAN) (Goodfellow et al. (2020)) in specific tasks within the field of computer vision. Diffusion models encompass a forward diffusion process and a reverse denoising process, each exhibiting Markovian behavior. The forward process incrementally introduces Gaussian noise into the original image, culminating in a transition to standard Gaussian noise. The forward diffusion process can be represented as $q(x_{t}|x_{t-1})=N(x_{t};\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I)$, where $t$ denotes the time step of each noise addition. The reverse denoising process employs the U-Net (Ronneberger et al. (2015)) architecture to accurately model the noise distribution at each timestep $t$. The image synthesis is achieved through a sequential denoising and sampling procedure, initiated from standard Gaussian noise.
In the context of image generation tasks, the conventional diffusion model executes two Markov processes in a large pixel space, resulting in substantial computational resource utilization. To address this issue, Latent Diffusion Models (LDM) (Rombach et al. (2022)) employs a VQ-VAE (Van Den Oord et al. (2017)) encoder to transform the pixel space into a low-dimensional latent space. Subsequently, the diffusion model’s training and generation are performed in the latent space, with the final generated image obtained by utilizing the VQ-VAE decoder. This approach significantly reduces computational resource requirements and inference time while preserving the quality of generated images.
@@ A.3 Diffusion Models for Video Generation
After achieving significant progress in text-to-image (T2I) generation tasks, diffusion models have piqued the interest of researchers in exploring their potential for text-to-video (T2V) generation. The pioneering work by Ho et al. (2022b) , introducing the 3D diffusion U-Net, marked significant progress in applying Diffusion Models to video generation. This was followed by further advancements by Ho et al. (2022a) , who utilized a cascaded sampling framework and super-resolution method to generate high-resolution videos. Subsequent contributions have expanded upon this work, notably with the incorporation of a temporal attention mechanism over frames by Singer et al. (2022) in Make-A-Video. Zhou et al. (2022) with MagicVideo, and He et al. (2022b) with LVDM, have integrated this mechanism into latent Diffusion Models, significantly enhancing video generation capabilities. However, due to the scarcity of paired text-video datasets and the high memory requirements for training 3D U-Nets, alternative approaches are being explored. These involve refining pre-trained T2I models to directly undertake T2V tasks. Khachatryan et al. (2023) introduced two enhancements to enable zero-shot adaptation of T2I models to T2V tasks: (1) the implementation of cross-frame attention, ensuring that the generation of each current frame in a video considers information from preceding frames; and (2) the consideration of inter-frame correlations during noise sampling, rather than random sampling for each frame independently. Wu et al. (2023) also employed cross-frame attention and achieved one-shot video editing by fine-tuning partial model parameters on individual videos.
@@
Appendix B Data Preprocessing
For the stimulus videos of the three datasets described below, we segmented them into 2-second clips, down-sampled the frame rate to 4Hz (i.e. evenly extracting 8 frames), then centrally cropped each frame, and resized each to a shape of 512$\times$512. Following the approach of Chen et al. (2024), we employed BLIP2 (Li et al. (2023)) to obtain textual descriptions for each video clip, with lengths not exceeding 20 words.
B.1 Video Captioning with BLIP2
Due to the absence of open-source video captioning models at the time of experimentation, we utilize the image captioning model BLIP2 to obtain text descriptions corresponding to each video clip. Two considerations are paramount in the design of the video captioning process: (1) the length of the text descriptions should not be excessively long, and (2) the text descriptions must reflect the scene transitions within the video segments. To achieve the first objective, we employ the following prompt: ’Question: What does this image describe? Answer in 20 words or less. Answer:’. Regarding the second point, we extracte 1st and 6th frames from every set of 8 frames, inputting them into BLIP2 to obtain their text descriptions. Subsequently, we calculate the CLIP similarity between each of the two text descriptions and 3rd frame. If the difference is no more than 0.05, it indicates minimal scene change, leading to the random selection of one text description from either 1st or 6th frame to represent the video clip. Otherwise, indicating a scene transition, we concatenate the two text descriptions with ’, then’ to provide a comprehensive textual description of the video clip. PyTorch code for the video captioning process is depicted in Algorithm 1.
@@ B.1 Video Captioning with BLIP2
1import torch
2import clip
3import random
4import numpy as np
5from PIL import Image
6from lavis.models import load_model_and_preprocess
7
8device = torch.device(’cuda:5’)
9clip_model, clip_preprocess = clip.load("ViT-B/32", device=device)
10model, vis_processors, _ = load_model_and_preprocess(name = "blip2_t5", model_type = "pretrain_flant5xxl", is_eval = True, device = device)
11prompt = "Question: What does this image describe? Answer in 20 words or less. Answer:"
12
13def Video_Captioning(Train_video_path_root):
14 Train_Captions = []
15 for i in tqdm(range(18)):
16 for j in tqdm(range(240)):
@@ 14 Train_Captions = []
18 frame1 = Image.open(frames_root + ’0000000.jpg’).convert("RGB")
19 frame1 = vis_processors["eval"](frame1).unsqueeze(0).to(device)
20 frame2 = Image.open(frames_root + ’0000056.jpg’).convert("RGB")
21 frame2 = vis_processors["eval"](frame2).unsqueeze(0).to(device)
22 frame_mid = Image.open(frames_root + ’0000024.jpg’).convert("RGB")
23 clip_image = clip_preprocess(frame_mid).unsqueeze(0).to(device)
24
25 caption1 = model.generate({"image": frame1, "prompt": prompt})
26 caption2 = model.generate({"image": frame2, "prompt": prompt})
27
28 text1 = clip.tokenize(caption1).to(device)
29 text2 = clip.tokenize(caption2).to(device)
30
31 with torch.no_grad():
32 image_features3 = clip_model.encode_image(clip_image)
33 text_features1 = clip_model.encode_text(text1)
34 text_features2 = clip_model.encode_text(text2)
35
36 cos_sim1 = torch.cosine_similarity(image_features3, text_features1)
37 cos_sim2 = torch.cosine_similarity(image_features3, text_features2)
38
39 if abs(cos_sim1 - cos_sim2) <= 0.05:
40 number = random.random()
41 if number >= 0.5:
42 caption = caption1[0]
43 else:
44 caption = caption2[0]
45 else:
46 caption = caption1[0] + ’, and then + caption2[0]
47
48 Train_Captions.append(caption)
49
50 Train_Captions = np.array(Train_Captions)
51 return Train_Captions
B.2 CC2017
@@
CC2017 (Wen et al. (2018)) dataset was first used in the work of Wen et al. (2018) This dataset include fMRI data from 3 subjects who view a variety of movie clips (23°$\times$23°) with a central fixation cross (0.8°$\times$0.8°). Clips are divided into 18 training movies and 5 testing movies, each eight minutes long, presented 2 and 10 times to each subject, respectively. MRI (T1 and T2-weighted) and fMRI data (2-second temporal resolution) are acquired using a 3-T system. The fMRI volumes are processed for artifact removal, motion correction (6 DOF), registered to MNI space, and projected onto cortical surfaces coregistered to a template.
@@
To extract voxels in the activated visual areas, we calculate the correlation of the time series for each voxel’s activation across 2 trials within the training set. Subsequently, we apply Fisher’s z-transform to the computed correlations, average the results across 18 sessions, and identify the most significant 4500 voxels using a paired t-test to form a mask. This mask is computed separately for each of the 3 subjects on their respective training sets and then applied to both training and test set data, with the selected voxels averaged across trials. Following the work of Nishimoto et al. (2011) and Han et al. (2019) , we utilize BOLD signals with a 4-second lag to represent the movie stimulus responses, thereby accounting for the hemodynamic response delay.
B.3 HCP
@@
This dataset is part of the Human Connectome Project (HCP) (Marcus et al. (2011)), encompassing BOLD (blood-oxygen-level dependent) responses from 158 subjects. For the subsequent experiments, three subjects (100610, 102816, and 104416) are randomly selected from this dataset. Data acquisition is performed using a 7T MRI scanner with a spatial resolution of 1.6 millimeters and a repetition time (TR) of 1 second. The utilized BOLD signals undergo standard HCP preprocessing procedures, which include correction for head motion and distortion, high-pass filtering, and removal of temporal artifacts via independent component analysis (ICA). The preprocessed BOLD responses are then registered to the MNI standard space.
@@
Due to the difficulty in directly acquiring fMRI data from multiple trials, we directly utilize the parcellation of the human cerebral cortex proposed by Glasser et al. (2016) to extract voxels within the activated visual cortex. The resulting ROIs we extract include: V1, V2, V3, hV4, PPA, FFA, LO, PHC, MT, MST, and TPOJ, totaling 5820 voxels. Following the work of Nishimoto et al. (2011) and Han et al. (2019) , we utilize BOLD signals with a 4-second lag to represent the movie stimulus responses, thereby accounting for the hemodynamic response delay. Given that no prior work has conducted video reconstruction experiments on these three subjects from the HCP dataset, except for Wang et al. (2022), we randomly shuffle all video segments and allocate 90% for the training set, with the remaining 10% reserved for the test set.
B.4 Algonauts2021